OpenStudy (anonymous):

PLEEEEEEEEEEEEEEEEEEEEEEEEEEEASE Check my answer for a medal? Express the complex number in trigonometric form. -2 + 2(square root of three)i I got 4(cos(5pi/6) + i sin(5pi/6))

OpenStudy (anonymous):

pls :((((

OpenStudy (anonymous):

The modulo of the complex number is correct

OpenStudy (anonymous):

okay so instead of 5pi/6 it should be 2pi/3? I'm not sure why though

OpenStudy (anonymous):

Correct, its 2pi/3

OpenStudy (anonymous):

I did arctan (2*sqrt(3) /2)

OpenStudy (anonymous):

To find the angle of a complex number, you apply the following formula: \[\arg(z) = \frac{ \Im(z) }{ \Re(z) }\]

OpenStudy (anonymous):

\[\arg(z) = \tan^{-1} \frac{ \Im(z) }{ \Re(z) }\]

OpenStudy (anonymous):

I missed the arctan in the previous formula.

OpenStudy (anonymous):

You missed the negative sign in your arctan.

OpenStudy (anonymous):

arctan (2sqrt(3) / -2)?

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

converted to radians it's (-pi/3) right? so where does the 2 come from?

OpenStudy (anonymous):

You will get \[-\frac{ \pi }{ 3 }\]

OpenStudy (anonymous):

Recall that you can add an angle by 2pi at any time, especially to make it positive.

OpenStudy (anonymous):

-pi/3 + 2pi = 5pi/3 -pi/3 - 2pi = -7pi/3??? I feel like I'm missing something obvious..:/

OpenStudy (anonymous):

Well, do recall that arctan, along with other inverse trig functions, is a multi-valued function.

OpenStudy (anonymous):

-pi/3 is only one of the possible solutions.

OpenStudy (anonymous):

Remember tangent is negative in both the second and fourth quadrants.

OpenStudy (anonymous):

I don't know how to find a value in another quadrant if it's not on the unit circle

OpenStudy (anonymous):

It's on the unit circle. You know that your first solution is -pi/3, which is 5pi/3 in the fourth quadrant.

OpenStudy (anonymous):

Your second solution must be \[\pi - \frac{ \pi }{ 3 } = \frac{ 2 \pi }{ 3 }\]

OpenStudy (anonymous):

why do you subtract it from pi?

OpenStudy (anonymous):

It's the unit circle. You subtract from the quadrant. You subtract from pi for the second quadrant.

OpenStudy (anonymous):

Oh wow. I never learned that... Thank you so much for your help!